40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5470     Entry time: Mon Sep 19 21:19:25 2011
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: Broadband photodiode characterization 

Another Hamasutu S3399 photodiode was tested with the electronic circuit as described in LIGO-D-1002969-v.

RF transimpedance is 1k although the DC transimpedance is 2k.

The noise level is 25pA/sqrt(Hz) which corresponds to a dark current of 1.9mA or 1.7mA in the independent measurement.

At all frequencies the noise is larger compared to Koji's measurement (see labbook page 4778).

 


In file idet_S3399.pdf the first point is not within its error bars on the fitted curved. This point corresponds to the dark noise measurement

I made this measurement again. Now it is on the fitted curve. In the previous measurement I pushed the save button a bit too early. The

averaging process has not been ready while I pushed the 'save'  button.

Dark current is 1.05mA and noise is lower than in the previous measurement.

New file are the XXX_v2.pdf files

current_noise_S3399_v2.pdf

 idet_S3399_v2.pdf

 


 idet_S3399.pdf

 current_noise_S3399.pdf

S3399_response.pdf

ELOG V3.1.3-